Circular No. 6737

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOTICE TO CONTRIBUTORS The following points can be made concerning the account on IAUC

6736. Firstly, some sources of information about minor planets are apparently unreliable. Secondly, apparent absence of motion over the course of less than several hours of an object near the ecliptic is not a guarantee that a minor planet has not been observed, particularly if it is near a stationary point. Thirdly, spectroscopic confirmation by professional astronomers using large telescopes is an expensive proposition and should be considered only as a final resort. Fourthly, the use of the WWW for confirmation purposes is potentially dangerous, in that if the Central Bureau is not cognizant of who knows of a purported discovery, a genuine discoverer could conceivably be deprived of credit.

In the specific case of a CCD or photographic detection of an apparent supernova and nova, observers are henceforth requested, before making a report, to perform at least three of the following four tasks, if credit is expected for a discovery. Firstly, they should make a precise astrometric measurement of the object’s position, specifying date and time (and magnitude); secondly, they should also observe on a second night to verify that the object is in precisely the same location; thirdly, they should show that the object was not present on comparable images on some other occasion; fourthly, they should clearly demonstrate the object’s nature from its spectrum. The first two points, in this case showing motion, are routinely expected of CCD and photographic discoverers of minor planets and comets. We recognize that the fourth point is difficult for all but experienced professional astronomers, but amateurs can still perform the first three. We also recognize that the first two points, and maybe even the third, represent a problem for visual discoverers, although a visual supernova hunter, in particular, can in this instance clearly profit by collaborating with a colleague who utilizes a CCD. If bad weather is expected, it may also be reasonable for a CCD observer to collaborate with a colleague elsewhere for the second-night follow-up, but this should be a trusted colleague who is contacted personally, not someone who responds to a mass e-mailing or WWW placing. The proliferation of internet communication greatly reduces the security involving requests for confirmation, particularly when the CCD images themselves are posted. This point was discussed at length at the meeting of IAU Commission 6 in Kyoto last month, in that such extensive use of the internet could completely jeopardize what is meant by a discovery and the proper assignment of credit. The matter is of particular import in the case of comets, because the IAU clearly wishes to continue the tradition of naming these objects for their discoverers.

1997 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden